OpenStudy (anonymous):

The cube of a number is the same as twice the square of the number. Find the number.

OpenStudy (anonymous):

Translate x^3 = 2x^2 Subtract 2x^2 from both sides x^3 - 2x^2 = 0 Factor x^2(x - 2) = 0 We have two solutions 0 and 2

OpenStudy (karatechopper):

i have noticed u have posted types of questions tht can be solved the same way. y dont u try one for urself?

OpenStudy (anonymous):

the number is 2 let x be the number x^3=2x^2 (2)^3=2(2)^2 8=8